OpenStudy (anonymous):

a) Find dy/dx given that x2 +y2 −4x+7y=15. (b) Under what conditions on x and/or y is the tangent line to this curve horizontal? Vertical? Help ?

OpenStudy (anonymous):

Have you done a) yet?

OpenStudy (anonymous):

@greenglasses I got stuck there :( I got 2x+2y+3 but the reference paper has it as (2x-4)+(2y+7)

OpenStudy (anonymous):

It's\[x^{2}+y^{2}-4x+7y=15\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you show me your work, I think you're making a similar mistake to last time.

OpenStudy (anonymous):

dy/dx(for all): x2=2x y2=2y -4x=-4 7y=7 2x+2y-4+7=0 (2x+2y)+3+0 (2x+2y)=-3 -3/(2x+2y)

OpenStudy (anonymous):

I know where you're going wrong. Has your teacher taught you implicit differentiation?

OpenStudy (anonymous):

I think that was the lesson I missed...

OpenStudy (anonymous):

Well, implicit differentiation tells you that \[(d/dx)(y) = dy/dx = y'\] right?

OpenStudy (anonymous):

So the derivative of any value of y to a power will be\[(d/dx)(y^{n})=ny^{n-1}y'\]

OpenStudy (anonymous):

Thanks a lot. I'm going to try to review this section

OpenStudy (anonymous):

Alright good luck :)

OpenStudy (anonymous):

Thanks for the help and patience! I'm trying to do some last minute cramming for a test

OpenStudy (anonymous):

Hey, I know how that feels, I'm doing that too here (for a different topic).

OpenStudy (anonymous):

Thanks for taking the time to help me. I wish you lots of luck on your test too! We can do it!!

OpenStudy (anonymous):

Yes, believing is half the battle ;)